# MV-VTON: Multi-View Virtual Try-On with Diffusion Models

### v2

@@ 3.3. Joint Attention Block
Specifically, in the $i$-th joint attention block, we first calculate self-attention for the current features $f_{in}^{i}$, followed by double cross-attention where the queries (Q) come from $f_{in}^{i}$ and global features $c_{g}$ serve as one set of keys (K) and values (V), while local features $c_{l}^{i}$ serve as another set of keys (K) and values (V).
After aligning to the person’s pose through cross-attention, the clothing features $c_{g}$ and $c_{l}^{i}$ are selectively fused in channel-wise dimension, i.e.,
(5) $f_{out}^{i}=softmax(\frac{Q^{i}_{g}(K^{i}_{g})^{T}}{\sqrt{d}})V^{i}_{g}+\lambda\odot softmax(\frac{Q^{i}_{l}(K^{i}_{l})^{T}}{\sqrt{d}})V^{i}_{l}~{},$
@@
where $Q^{i}_{g},K^{i}_{g},V^{i}_{g}$ represent the Q, K, V of global branch, $Q^{i}_{l},K^{i}_{l},V^{i}_{l}$ represent the Q, K, V of local branch, $\lambda$ is the learnable fusion vector, $\odot$ represents channel-wise multiplication, and $f_{out}^{i}$ represents the clothing features after selective fusion. By engaging and fusing the global and local features of garments, we can enhance the retention of high-frequency details like texts and patterns in original clothing.
3.4. Training Objectives
@@
As stated in Section 3.1, diffusion models learn to generate images from random Gaussian noise. However, the training objective in Eq. (3) is performed in latent space, and does not explicitly constrain the generated results in visible image space, resulting in slight differences in color from the ground truth. To alleviate the problem, we additionally employ $\ell_{1}$ loss $\mathcal{L}_{1}$ and perceptual loss (Johnson et al., 2016) $\mathcal{L}_{perc}$. The $\mathcal{L}_{1}$ loss is calculated by
(6) $\mathcal{L}_{1}=\left\|\hat{x}-x\right\|_{1},$
@@
(7) $\mathcal{L}_{perc}=\sum_{k=1}^{5}\left\|\phi_{k}(\hat{x})-\phi_{k}(x)\right\|_{1},$
where $\phi_{k}$ represents the $k$-th layer of VGG (Simonyan and Zisserman, 2014). Totally, the overall training objective can be written as,
(8) $\mathcal{L}=\mathcal{L}_{LDM}+\lambda_{1}\mathcal{L}_{1}+\lambda_{perc}\mathcal{L}_{perc}~{},$
@@
Table 1. Quantitative comparison with previous work on paired setting. For multi-view virtual try-on task, we show results on our proposed MVG dataset. For frontal-view virtual try-on task, we show results on VITON-HD dataset (Choi et al., 2021) and DressCode dataset (Morelli et al., 2022). The best results have been bolded. Note that all previous works have been finetuned on our proposed MVG dataset when comparing on multi-view virtual try-on task.
Method MVG VITON-HD DressCode - Upper Body
@@
Paint by Example (Yang et al., 2023) 0.120 0.880 54.38 14.95 0.150 0.843 13.78 4.48 0.078 0.899 15.21 4.51
PF-AFN (Ge et al., 2021b) 0.139 0.873 49.47 12.81 0.141 0.855 7.76 4.19 0.091 0.902 13.11 6.29
GP-VTON (Xie et al., 2023) - - - - 0.085 0.889 6.25 0.77 0.236 0.781 19.37 8.07
LaDI-VTON (Morelli et al., 2023) 0.069 0.921 29.14 4.39 0.094 0.872 7.08 1.49 0.063 0.922 11.85 3.20
DCI-VTON (Gou et al., 2023) 0.062 0.929 25.71 0.95 0.074 0.893 5.52 0.57 0.043 0.937 11.87 1.91
StableVITON (Kim et al., 2023) 0.063 0.929 23.52 0.46 0.073 0.888 6.15 1.34 0.040 0.937 10.18 1.70
Ours 0.050 0.936 22.17 0.36 0.068 0.897 5.42 0.48 0.040 0.941 8.26 1.39
Table 2. Unpaired setting’s quantitative results on our MVG dataset and VITON-HD dataset (Choi et al., 2021). The best results have been bolded.
@@
Paint by Example (Yang et al., 2023) 43.79 5.92 17.27 4.56
PF-AFN (Ge et al., 2021b) 47.38 7.04 21.18 6.57
GP-VTON (Xie et al., 2023) - - 9.11 1.21
LaDI-VTON (Morelli et al., 2023) 36.61 3.39 9.55 1.83
DCI-VTON (Gou et al., 2023) 36.03 3.79 8.93 1.07
StableVITON (Kim et al., 2023) 35.85 4.22 9.86 1.09
Ours 34.48 2.92 8.67 0.78
4. experiments
4.1. Experiments Setting
Multi-View Dataset: For the proposed multi-view virtual try-on task, we collect MVG dataset containing 1,009 samples. Each sample contains five images of the same person wearing the same garment from five different views, for a total of 5,045 images, as shown in Figure 2(b). The image resolution is about 1K. Collection details of this dataset are provided in the supplementary material. In this work, we split it into a training set with 812 samples and a test set with 197 samples. During training, we extract the clothing from the frontal and back view images (the first and fifth columns of Figure 2(b)) to reconstruct the try-on results from other views (the second to fourth columns of Figure 2(b)), and use the real images from the corresponding views as ground truth. During testing, users only need to provide two images in frontal and back views to generate try-on results of a person from multiple views.
Frontal-View Datasets. The proposed method can also be applied to frontal-view virtual try-on task. Our frontal-view experiments are carried out on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. VITON-HD (Choi et al., 2021) has 13,679 frontal-view person and upper-body clothing image pairs with a resolution of 1,024×768. DressCode (Morelli et al., 2022) contains three categories: upper-body clothing, lower-body clothing, and dresses, with a total of more than 53,000 image pairs of frontal-view clothing and persons. Among them, upper-body clothing contains 15,363 samples. For the division of the datasets, we follow the previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023). For VITON-HD (Choi et al., 2021) dataset, we use 2,032 pairs as the test set and 11,647 pairs as the training set. For DressCode (Morelli et al., 2022) dataset, we employ 1,800 pairs as the test set, and the rest as the training set.
Evaluation Metrics. Following previous works (Kim et al., 2023; Morelli et al., 2023; Gou et al., 2023; Ge et al., 2021b; Xie et al., 2023; Yang et al., 2023), we use four metrics to evaluate the performance of our method: Structural Similarity (SSIM) (Wang et al., 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018), Frechet Inception Distance (FID) (Heusel et al., 2017) and Kernel Inception Distance (KID) (Bińkowski et al., 2018). LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004) tend to measure the consistency of the generated try-on results with ground truth. FID (Heusel et al., 2017) and KID (Bińkowski et al., 2018) are employed to measure the realism of images. Specifically, for paired test setting, which means directly using the paired data in the dataset, we utilize the above four metrics for evaluation. For unpaired test setting, which means that the given garment is different from the garment originally worn by target person, we use FID and KID for evaluation, and in order to distinguish them from the paired setting, we named them ${\rm FID_{u}}$ and ${\rm KID_{u}}$ respectively.
Figure 5. Qualitative comparison on multi-view virtual try-on task.
[…]
[Figure 5. Fully described in the text.]
Figure 6. Qualitative comparison on frontal-view virtual try-on task.
[…]
[Figure 6. Fully described in the text.]
Implementation Details. We use Paint by Example (Yang et al., 2023) as the backbone of our method and copy the weights of its encoder to initialize $\mathcal{E}_{l}$. The hyper-parameter $\lambda_{1}$ is set to 1e-1, and $\lambda_{perc}$ is set to 1e-4. We train our model on 2 NVIDIA Tesla A100 GPUs for 40 epochs with a batch size of 4 and a learning rate of 1e-5. We use AdamW (Loshchilov and Hutter, 2017) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$.
Comparison Settings. We compare our method with Paint By Example (Yang et al., 2023), PF-AFN (Ge et al., 2021b), GP-VTON (Xie et al., 2023), LaDI-VTON (Morelli et al., 2023), DCI-VTON (Gou et al., 2023) and StableVITON (Kim et al., 2023) on both frontal-view and multi-view virtual try-on tasks. For multi-view virtual try-on, we compare these methods on the proposed MVG dataset. For the sake of fairness, we fine-tune the previous methods on the MVG dataset according to its original training settings. Since previous methods can only input a single clothing image, we input frontal and back view clothing respectively and select the best result. For frontal-view virtual try-on, we compare these methods on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. Following previous works’ settings, the proposed MV-VTON only inputs one frontal-view garment during training and inference. When evaluating previous methods, we directly use their released models to test, and train the models if not available.
4.2. Quantitative Evaluation
Table 1 reports the quantitative results on the paired setting, and Table 2 shows the unpaired setting’s results. On the multi-view virtual try-on task, as can be seen, thanks to the view-adaptive selection mechanism, our method can reasonably select clothing features according to the person’s pose, so it is better than existing methods in various metrics, especially on LPIPS and SSIM. Furthermore, owing to the joint attention block, our approach excels in preserving high-frequency details of the original garments across both frontal-view and multi-view virtual try-on scenarios, thus achieving superior performance in these metrics.
4.3. Qualitative Evaluation
Multi-View Virtual Try-On. As shown in Figure 5, MV-VTON generates more realistic multi-view results compared to the previous five methods. Specifically, in the first row of Figure 5, due to the lack of adaptive selection of clothes, previous methods have difficulty in generating hoods of the original cloth. Moreover, in the second and third rows, previous methods often struggle to maintain fidelity to the original garments. In contrast, our method effectively addresses the aforementioned problems and generates high-fidelity results.
Frontal-View Virtual Try-On. As shown in Figure 6, our method also demonstrates superior performance over existing methods on frontal-view virtual try-on task, particularly in retaining clothing details. Specifically, in the second row of the Figure 6, our method better preserves the literal ’KORS’ in the clothing. In addition, benefiting from the selective fusion of global and local clothing features in joint attention blocks, our method not only preserves the shape of special garments (in the first row), but also faithfully generates complex patterns (in the third row).
Figure 7. Visual comparison of effect of the proposed view-adaptive selection mechanism on MVG dataset.
[…]
[Figure 7. Fully described in the text.]
Figure 8. Visual comparison of effect of the proposed joint attention block on VITON-HD (Choi et al., 2021) dataset.
[…]
[Figure 8. Fully described in the text.]
4.4. User Study
We conduct a user study involving 28 participants, each of which is presented with test results generated by our and previous methods. The participants evaluate the results based on fidelity (indicating whether the generated clothing is faithful to original clothing) and identity (indicating whether human body details such as the face and hands are preserved well) and choose the best one among all methods. We then calculate the proportion of the best images to the total results for each method. As shown in Table 3, the proposed method outperforms previous methods both on fidelity and identity.
4.5. Ablation Study
Effect of View-Adaptive Selection. We investigate the effect of view-adaptive selection on the multi-view virtual try-on task. Specifically, on the basis of no clothing selection, we add hard-selection when extracting global features and soft-selection when extracting local features, respectively. No hard-selection represents that we directly concatenate two garments’ features encoded by CLIP (Radford et al., 2021), and no soft-selection means that two clothing features are concatenated without passing soft-selection blocks. Comparison results are shown in Table 4 and Figure 7. As can be seen, without any selection, the performance is greatly reduced. Merely employing hard-selection or soft-selection can enhance performance, but the generated results still exhibit some unnaturalness, such as the unrealistic hoods and black stripes, as shown in Figure 7. In contrast, by combining hard-selection and soft-selection, our method can generate more faithful try-on results.
Table 3. User study results. We make comparisons with previous works on MVG and VITON-HD (Choi et al., 2021) datasets.
Method MVG VITON-HD
Fidelity↑ Identity↑ Fidelity↑ Identity↑
Paint by Example (Yang et al., 2023) 0.76% 0.33% 0.32% 1.66%
PF-AFN (Ge et al., 2021b) 0.37% 0.63% 0.17% 0.19%
GP-VTON (Xie et al., 2023) - - 11.4% 0.42%
LaDI-VTON (Morelli et al., 2023) 11.38% 19.81% 3.37% 9.57%
DCI-VTON (Gou et al., 2023) 3.77% 8.3% 14.13% 13.63%
StableVITON (Kim et al., 2023) 15.13% 29.2% 17.9% 27.42%
Ours 68.59% 41.73% 52.71% 47.11%
Table 4. Ablation study of our proposed view-adaptive selection mechanism on MVG dataset.
@@ 4.5. Ablation Study
Table 5. Ablation study of the joint attention block on MVG and VITON-HD (Choi et al., 2021) datasets.
@@
Effect of Joint Attention Block. In order to demonstrate the effectiveness of fusing global and local features through joint attention blocks, we discard the global feature extraction branch and the local feature extraction branch respectively. Results are shown in Table 5 and Figure 8. As can be seen, relying solely on global features may lead to loss of details, such as the distorted text ’VANS’ in the first row and the missing letter ’C’ in the second row. Moreover, if only local features are provided, the results may also have unfaithful textures, such as artifacts on the person’s chest. Compared to them, we fuse global and local features through joint attention blocks, which can refine details in garments while preserving semantic information.
5. conclusion
@@ Appendix A IMPLEMENTATION DETAILS
Hard-Selection. In this section, we present more details about the proposed hard-selection for global clothing features. Specifically, in multi-view virtual try-on task, we use OpenPose (Cao et al., 2017; Simon et al., 2017; Wei et al., 2016) to extract the skeleton images of target person, frontal clothing and back clothing as pose information $p_{h}$, $p_{f}$, and $p_{b}$, respectively. After that, we decide whether to use frontal-view clothing or back-view clothing based on the relative positions of the target person’s left arm and right arm in the skeleton images. As shown in Figure A, if the right arm appears positioned to the left of the left arm in the skeleton image (columns one to three in Figure A), frontal-view clothing is chosen; otherwise, back-view clothing is preferred (columns four to five in Figure A). In addition, following previous works (Gou et al., 2023; Xie et al., 2023; Morelli et al., 2023), we adopt PF-AFN (Ge et al., 2021b) and GP-VTON (Xie et al., 2023) to obtain the pre-warped cloth on frontal-view try-on task; and we utilize StableVITON (Kim et al., 2023) to get it on multi-view try-on task.
Pose Encoder. The pose encoder is used to extract features of skeleton images. It is a tiny network that contains three blocks, followed by the layer normalization (Ba et al., 2016). Each block comprises one convolution layer, one GELU (Hendrycks and Gimpel, 2016) activation layer, and a down-sampling operation. We utilize the acquired pose embeddings as input for the proposed soft-selection block.
MVG Dataset. To construct dataset for multi-view virtual try-on (MV-VTON), we first collect a large number of videos from YOOX NET-A-PORTER [^1] https://net-a-porter.com, Taobao [^2] https://taobao.com and TikTok [^3] https://douyin.com websites, then filter out 1009 videos where the person in the video turns at least 180 degrees while wearing clothes. Afterwards, we divide each video into frames and handpick 5 frames to constitute a sample within our MVG dataset. Across these 5 frames, the person is captured from various angles, approximately spanning 0 (i.e., frontal view), 45, 90, 135, and 180 (i.e., back view) degrees, as shown in the first row of Figure B. In addition, following the previous DressCode dataset (Morelli et al., 2022), we employ SCHP model (Li et al., 2020) to extract the corresponding human parsing maps and utilize DensePose (Güler et al., 2018) to obtain the person’s dense labels, as shown in the second and third row of Figure B. Human parsing maps can be utilized to generate cloth-agnostic person images, which are necessary for training and inference processes.

### v4

@@ Introduction
Specifically, in the $i$-th joint attention block, we first calculate self-attention for the current features $f_{in}^{i}$. Then, we deploy a double cross-attention. The queries (Q) come from $f_{in}^{i}$ and global features $c_{g}$ serve as one set of keys (K) and values (V), while local features $c_{l}^{i}$ serve as another set of keys (K) and values (V). After aligning to the person’s pose through cross-attention, the clothing features $c_{g}$ and $c_{l}^{i}$ are selectively fused in channel-wise dimension, i.e.,
$f_{out}^{i}={}$ $softmax(\frac{Q^{i}_{g}(K^{i}_{g})^{T}}{\sqrt{d}})V^{i}_{g}~{}+$ (5)
$\lambda\odot softmax(\frac{Q^{i}_{l}(K^{i}_{l})^{T}}{\sqrt{d}})V^{i}_{l}~{},$
@@
where $Q^{i}_{g},K^{i}_{g},V^{i}_{g}$ represent the Q, K, V of global branch, $Q^{i}_{l},K^{i}_{l},V^{i}_{l}$ represent the Q, K, V of local branch, $\lambda$ is the learnable fusion vector, $\odot$ represents channel-wise multiplication, and $f_{out}^{i}$ represents the clothing features after selective fusion. By engaging and fusing the global and local clothing features, we can enhance the retention of high-frequency garment details, e.g., texts and patterns.
Training Objectives
@@
As stated in preliminaries, diffusion models learn to generate images from random Gaussian noise. However, the training objective in Eq. (3) is performed in latent space, and does not explicitly constrain the generated results in visible image space, resulting in slight differences in color from the ground truth. To alleviate the problem, we additionally employ $\ell_{1}$ loss $\mathcal{L}_{1}$ and perceptual loss (Johnson, Alahi, and Fei-Fei 2016) $\mathcal{L}_{perc}$. The $\mathcal{L}_{1}$ loss is calculated by
$\mathcal{L}_{1}=\left\|\hat{x}-x\right\|_{1},$ (6)
@@
$\mathcal{L}_{perc}=\sum_{k=1}^{5}\left\|\phi_{k}(\hat{x})-\phi_{k}(x)\right\|_{1},$ (7)
where $\phi_{k}$ represents the $k$-th layer of VGG (Simonyan and Zisserman 2014). Totally, the overall training objective can be written as,
$\mathcal{L}=\mathcal{L}_{LDM}+\lambda_{1}\mathcal{L}_{1}+\lambda_{perc}\mathcal{L}_{perc}~{},$ (8)
@@
Experiments
Experiments Settings
Datasets: For the proposed multi-view virtual try-on task, we collect MVG dataset containing 1,009 samples. Each sample contains five images of the same person wearing the same garment from five different views, for a total of 5,045 images, as shown in Figure 2(b). The image resolution is about 1K. We’ll explain how the datasets are collected and how they’re used for MV-VTON in the supplementary material. The proposed method can also be applied to frontal-view virtual try-on task. Our frontal-view experiments are carried out on VITON-HD (Lee et al. 2022) and DressCode (Morelli et al. 2022) datasets. They contain more than 10,000 frontal-view person and upper-body clothing image pairs. We follow previous work for the use of them.
Evaluation Metrics. Following previous works (Kim et al. 2023; Morelli et al. 2023), we use four metrics to evaluate the performance of our method: Structural Similarity (SSIM) (Wang et al. 2004), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. 2018), Frechet Inception Distance (FID) (Heusel et al. 2017) and Kernel Inception Distance (KID) (Bińkowski et al. 2018). Specifically, for paired test setting, which means directly using the paired data in the dataset, we utilize the above four metrics for evaluation. For unpaired test setting, which means that the given garment is different from the garment originally worn by target person, we use FID and KID for evaluation, and in order to distinguish them from the paired setting, we named them ${\rm FID_{u}}$ and ${\rm KID_{u}}$ respectively.
Methods Reference MVG VITON-HD DressCode - Upper Body
@@
Paint by Example CVPR23 0.120 0.880 54.38 14.95 0.150 0.843 13.78 4.48 0.078 0.899 15.21 4.51
PF-AFN CVPR21 0.139 0.873 49.47 12.81 0.141 0.855 7.76 4.19 0.091 0.902 13.11 6.29
GP-VTON CVPR23 - - - - 0.085 0.889 6.25 0.77 0.236 0.781 19.37 8.07
LaDI-VTON MM23 0.069 0.921 29.14 4.39 0.094 0.872 7.08 1.49 0.063 0.922 11.85 3.20
DCI-VTON MM23 0.062 0.929 25.71 0.95 0.074 0.893 5.52 0.57 0.043 0.937 11.87 1.91
StableVITON CVPR24 0.063 0.929 23.52 0.46 0.073 0.888 6.15 1.34 0.040 0.937 10.18 1.70
IDM-VTON ECCV24 0.095 0.896 34.66 5.33 0.135 0.826 14.36 8.63 0.066 0.912 13.88 5.39
Ours - 0.050 0.936 22.18 0.35 0.069 0.897 5.43 0.49 0.040 0.941 8.26 1.39
Table 1: Quantitative comparison with previous work on paired setting. For multi-view virtual try-on task, we show results on our proposed MVG dataset. For frontal-view virtual try-on task, we show results on VITON-HD dataset (Lee et al. 2022) and DressCode dataset (Morelli et al. 2022). The best results have been bolded. Note that all previous works have been finetuned on our proposed MVG dataset when comparing on multi-view virtual try-on task.
Figure 5: Qualitative comparisons on multi-view virtual try-on task with MVG dataset.
Figure 6: Qualitative comparisons on frontal-view virtual try-on task with VITON-HD and DressCode datasets.
Implementation Details. We use Paint by Example (Yang et al. 2023) as the backbone of our method and copy the weights of its encoder to initialize $\mathcal{E}_{l}$. The hyper-parameter $\lambda_{1}$ is set to 1e-1, and $\lambda_{perc}$ is set to 1e-4. We train our model on 2 NVIDIA Tesla A100 GPUs for 40 epochs with a batch size of 4 and a learning rate of 1e-5. We use AdamW (Loshchilov and Hutter 2017) optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$.
Comparison Settings. We compare our method with Paint By Example (Yang et al. 2023), PF-AFN (Ge et al. 2021b), GP-VTON (Xie et al. 2023), LaDI-VTON (Morelli et al. 2023), DCI-VTON (Gou et al. 2023), StableVITON (Kim et al. 2023) and IDM-VTON (Choi et al. 2024) on both frontal-view and multi-view virtual try-on tasks. For multi-view virtual try-on, we compare these methods on the proposed MVG dataset. For the sake of fairness, we fine-tune the previous methods on the MVG dataset according to its original training settings. Since previous methods can only input a single clothing image, we input frontal and back view clothing respectively and select the best result. For frontal-view virtual try-on, we compare these methods on VITON-HD and DressCode datasets. Following previous works’ settings, the proposed MV-VTON only inputs one frontal-view garment during training and inference.
Quantitative Evaluation
Table 1 reports the quantitative results on the paired setting, and Table 2 shows the unpaired setting’s results. On the multi-view virtual try-on task, as can be seen, thanks to the view-adaptive selection mechanism, our method can reasonably select clothing features according to the person’s pose, so it is better than existing methods in various metrics, especially on LPIPS and SSIM. Furthermore, owing to joint attention blocks, our approach excels in preserving high-frequency details of the original garments across both frontal-view and multi-view virtual try-on scenarios, thus achieving superior performance in these metrics.
Qualitative Evaluation
Multi-View Virtual Try-On. As shown in Figure 6, MV-VTON generates more realistic multi-view results compared to the previous five methods. Specifically, in the first row, due to the lack of adaptive selection of clothes, previous methods have difficulty in generating hoods of the original cloth. Moreover, in the second row, previous methods often struggle to maintain fidelity to the original garments. In contrast, our method effectively addresses the aforementioned problems and generates high-fidelity results. We provide more results of multi-view virtual try-on in the supplementary materials.
@@
Paint by Example 43.79 5.92 17.27 4.56
PF-AFN 47.38 7.04 21.18 6.57
GP-VTON - - 9.11 1.21
LaDI-VTON 36.61 3.39 9.55 1.83
DCI-VTON 36.03 3.79 8.93 1.07
StableVITON 35.85 4.22 9.86 1.09
IDM-VTON 40.73 5.74 18.27 10.43
Ours 33.44 2.69 8.67 0.78
Table 2: Unpaired setting’s quantitative results on our MVG dataset and VITON-HD dataset. The best results have been bolded.
@@
Table 3: Ablation study of our proposed view-adaptive selection mechanism on MVG dataset.
@@
Table 4: Ablation study of joint attention blocks on MVG and VITON-HD datasets.
Frontal-View Virtual Try-On. As shown in Figure 6, our method also demonstrates superior performance over existing methods on frontal-view virtual try-on task, particularly in retaining clothing details. Specifically, our method not only faithfully generates complex patterns (in the first row), but also better preserves the literal ’Wrangler’ in the clothing (in the second row). We provide more qualitative comparisons in the supplementary materials, as well as dressing results under complex human pose conditions.
Ablation Studies
Effect of View-Adaptive Selection. We investigate the effect of view-adaptive selection on the multi-view virtual try-on task. Specifically, no hard-selection represents that we directly concatenate two garments’ features encoded by CLIP, and no soft-selection means that two clothing features are concatenated without passing soft-selection blocks. Comparison results are shown in Table 3 and Figure 7. As can be seen, the performance is greatly reduced without hard-selection and soft-selection. No hard-selection will confuse two view’s cloth features, as shown by the blurriness of the ’POP’ text in Figure 7. In addition, no soft-selection causes the model to lose some cloth information when processing the side view situation, such as the missing white hood and cuffs in Figure 7.
Figure 7: Visualization of view-adaptive selection’s effect.
Figure 8: Visualization of joint attention blocks’ effect.
Effect of Joint Attention Blocks. In order to demonstrate the effectiveness of fusing global and local features through joint attention blocks, we discard the global feature extraction branch and the local feature extraction branch respectively. Results are shown in Table 4 and Figure 8. As can be seen, relying solely on global features may lead to loss of details, such as the distorted text ’VANS’ in the first row and the missing letter ’C’ in the second row. Moreover, if only local features are provided, the results may also have unfaithful textures, such as artifacts on the person’s chest. Compared to them, we fuse global and local features through joint attention blocks, which can refine details in garments while preserving semantic information.
Conclusion
@@
Acknowledgments
This work was supported by the National Key R&D Program of China (2022YFA1004100).
@@ Appendix A IMPLEMENTATION DETAILS
Hard-Selection. In this section, we present more details about the proposed hard-selection for global clothing features. Specifically, in multi-view virtual try-on task, we use OpenPose (Cao et al. 2017; Simon et al. 2017; Wei et al. 2016) to extract the skeleton images of target person, frontal clothing and back clothing as pose information $p_{h}$, $p_{f}$, and $p_{b}$, respectively. After that, we decide whether to use frontal-view clothing or back-view clothing based on the relative positions of the target person’s left arm and right arm in the skeleton images. As shown in Figure A, if the right arm appears positioned to the left of the left arm in the skeleton image (columns one to three in Figure A), frontal-view clothing is chosen; otherwise, back-view clothing is preferred (columns four to five in Figure A). In addition, following previous works (Gou et al. 2023; Xie et al. 2023; Morelli et al. 2023), we adopt an additional warping network (Ge et al. 2021b; Kim et al. 2023) to obtain the pre-warped cloth.
Pose Encoder. The pose encoder is used to extract features of skeleton images. It is a tiny network that contains three blocks, followed by the layer normalization (Ba, Kiros, and Hinton 2016). Each block comprises one convolution layer, one GELU (Hendrycks and Gimpel 2016) activation layer, and a down-sampling operation. We utilize the acquired pose embeddings as input for the proposed soft-selection block.
MVG Dataset. To construct dataset for multi-view virtual try-on (MV-VTON), we first collect a large number of videos from YOOX NET-A-PORTER [^1] https://net-a-porter.com, Taobao [^2] https://taobao.com and TikTok [^3] https://douyin.com websites, then filter out 1009 videos where the person in the video turns at least 180 degrees while wearing clothes. Afterwards, we divide each video into frames and handpick 5 frames to constitute a sample within our MVG dataset. Across these 5 frames, the person is captured from various angles, approximately spanning 0 (i.e., frontal view), 45, 90, 135, and 180 (i.e., back view) degrees, as shown in the first row of Figure B. In addition, following the previous DressCode dataset (Morelli et al. 2022), we employ SCHP model (Li et al. 2020) to extract the corresponding human parsing maps and utilize DensePose (Güler, Neverova, and Kokkinos 2018) to obtain the person’s dense labels, as shown in the second and third row of Figure B. Human parsing maps can be utilized to generate cloth-agnostic person images, which are necessary for training and inference processes.
Figure A: Visualization of the person and corresponding poses. We select one of garments based on the relative positions of left and right arms in the skeleton image when performing hard-selection on the multi-view virtual try-on task.
Figure B: Examples of human parsing maps and dense pose in our dataset. The parsing maps can be used to synthesize cloth-agnostic person images.